# A multifluid model with chemically reacting components – construction of weak solutions

### v2

@@ 1 Description of the multifluid system
$\partial_{t}\varrho_{i}+\mathrm{div\,}(\varrho_{i}u)-\mathrm{div\,}F_{i}$ $=\omega_{i}(\vec{\varrho}\,),\quad i=1,\dots,N,$ in $\mbox{in \ }\Omega\times(0,T),$ (1.1)
$-\mu\Delta u-\nabla((\mu+\lambda)\mathrm{div\,}u)+\nabla p(\vec{\varrho}\,)$ $=0,$ in $\mbox{in \ }\Omega\times(0,T),$
@@
$u\cdot n$ $=0$ at $\mbox{at \ }\partial\Omega\times(0,T),$ (1.2)
$n\cdot\mathbb{T}(u,p)\cdot\tau_{k}+fu\cdot\tau_{k}$ $=0,\quad k=1,\dots,2,$ at $\mbox{at \ }\partial\Omega\times(0,T),$
$\quad F_{i}\cdot n$ $=0,\quad i=1,...,N,$ at $\mbox{at \ }\partial\Omega\times(0,T),$
@@
$p(\vec{\varrho}\,)=\sum_{i=1}^{N}p_{i}(\varrho_{i}),\mbox{ \ with \ }p_{i}(\varrho_{i}\,)=\frac{1}{m_{i}}\varrho_{i}^{\gamma_{i}},$ (1.3)
@@ 1.1 Main results
$12\gamma_{\max}<3\gamma_{\min}-\gamma_{S}+$ 1 (1.8)
@@
in $[0,T]\times\Omega$ with
[…]
@@
where $\vec{q}=(q_{1},\dots,q_{N})$ is a certain projection of $\left(\frac{\gamma_{1}}{\gamma_{1}-1}\frac{p_{1}}{\varrho_{1}},\dots,\frac{\gamma_{N}}{\gamma_{N}-1}\frac{p_{N}}{\varrho_{N}}\right)$ on the $N-1$-dimensional space orthogonal to $(1,\dots,1)$ (in particular $q_{N}=-\sum_{j=1}^{N-1}q_{j}$). It is shown that in the case when $\varrho_{1},\dots,\varrho_{N}\neq 0$, the coefficients $b_{i,j}$ from the relation (1.10) form an invertible matrix, which in turn provides an $L^{2}$ bound on $\nabla q_{i}$. Similar approach, with a different definition of the fluxes, was considered for example in [9, 11].
@@ 2 A priori estimates
$-2\sum_{i=1}^{N}\frac{\gamma_{i}}{m_{i}}\int_{\Omega}\varrho_{i}^{\gamma_{i}-1}\omega_{i}\;\mathrm{d}x\geq-2\frac{\gamma_{3}}{m_{3}}\int_{\Omega}\varrho_{3}^{\gamma_{3}-1}\omega_{3}\;\mathrm{d}x.$
@@
$\begin{array}[]{ll}-\mu\Delta\alpha=0&\mbox{in }\Omega,\\ \alpha\cdot\tau_{1}=(2\chi_{1}-f/\mu)u\cdot\tau_{1},\quad\alpha\cdot\tau_{2}=(f/\mu-2\chi_{2})u\cdot\tau_{2}&\mbox{at }\partial\Omega,\\ \frac{\partial(\alpha\cdot n)}{\partial n}=-(\alpha\cdot\tau_{1})_{\tau_{1}}-(\alpha\cdot\tau_{2})_{\tau_{2}}&\mbox{at }\partial\Omega.\\ \end{array}$ (2.6)
@@ 3.1 The Galerkin approximation
The equation (3.3) admits a global weak solution in $[0,T]\times\Omega$.
@@ 3.3 Boundedness of $\varrho$
$-2\sum_{i=1}^{N}\frac{1}{m_{i}}\int_{\Omega}\varrho_{i}\omega_{i}^{\delta}\;\mathrm{d}x\geq-2\frac{1}{m_{3}}\int_{\Omega}\varrho_{3}\omega_{3}^{\delta}\;\mathrm{d}x.$
@@
By performing elementary operations on the matrix $B=(b_{i,j})_{i,j=1,\dots,N-1}$, it is easy to see that $\det B=\frac{1}{\varrho}\varrho_{1}\dots\varrho_{N}$. Indeed, adding the verses $12,\dots,N-$ 1 to the first one and dividing it by $\frac{\varrho_{N}}{\varrho}$, we get the matrix
@@
$\det B=\frac{\varrho_{N}}{\varrho}\cdot\varrho_{1}\dots\varrho_{N-1}\text{ \ \ as we claimed.}$.
@@
[…]
then $\{\varrho_{k}\}_{k\in\mathbb{N}}$ is compact in $L^{p}((0,T)\times\Omega)$. Conversely, if $\{\varrho_{k}\}_{k\in\mathbb{N}}$ is compact in $L^{p}((0,T)\times\Omega)$, then the above quantity converges to 00.
@@
The last term has a good sign, so we just estimate it by 00. The first term is estimated in the same way as in [7]. By the definition of $K_{h}$, we have
@@
Denoting by $Ed$ an $H^{1}(\mathbb{R}^{3})$ extension of $d$, we have
@@
applied to $M|\nabla u^{x}|$ and $\int_{\Omega}\overline{K}_{h}(x-y)(\varrho^{x}-\varrho^{y})^{2}\mathrm{d}y$ extended by 00 outside $\Omega$. We get
@@ 5 The general diffusing/non-diffusing case
and computing $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{R}_{h}$, it is clear that for the non-diffusing components the equivalents of the terms $A_{1}$, $A_{2}$, $A_{4}$ and $A_{5}$ can be dealt with in the same manner, whereas $A_{3}$ is just equal to 00. In consequence, the computations from Section 4 give the strong convergence of densities for both diffusing and non-diffusing components.

### v3

@@ 1 Description of the multifluid system
$\partial_{t}\varrho_{i}+\mathrm{div\,}(\varrho_{i}u)-\mathrm{div\,}F_{i}$ $=\omega_{i}(\vec{\varrho}\,),\quad i=1,\dots,N,$ $\mbox{in \ }\Omega\times(0,T),$ (1.1)
$-\mu\Delta u-\nabla((\mu+\lambda)\mathrm{div\,}u)+\nabla p(\vec{\varrho}\,)$ $=0,$ $\mbox{in \ }\Omega\times(0,T),$
@@
$u\cdot n$ $=0$ $\mbox{at \ }\partial\Omega\times(0,T),$ (1.2)
$n\cdot\mathbb{T}(u,p)\cdot\tau_{k}+fu\cdot\tau_{k}$ $=0,\quad k=1,\dots,2,$ $\mbox{at \ }\partial\Omega\times(0,T),$
$\quad F_{i}\cdot n$ $=0,\quad i=1,.,N,$ $\mbox{at \ }\partial\Omega\times(0,T),$
@@
[…]
@@ 1.1 Main results
$2\gamma_{\max}<3\gamma_{\min}-\gamma_{S}+1$ (1.8)
@@
in $0,T]\times\Omega$ with
${\varrho_{i}}_{|_{t=0}}=\varrho_{0,i,\varepsilon}\in C^{\infty}(\Omega),\quad\varrho_{0,i,\varepsilon}\to\varrho_{0,i}\quad\text{in}\quad L^{2}(\Omega),$
@@
where $\vec{q}=(q_{1},\dots,q_{N})$ is a certain projection of $\left(\frac{\gamma_{1}}{\gamma_{1}-1}\frac{p_{1}}{\varrho_{1}},\dots,\frac{\gamma_{N}}{\gamma_{N}-1}\frac{p_{N}}{\varrho_{N}}\right)$ on the N1N-1-dimensional space orthogonal to $(1,\dots,1)$ (in particular $q_{N}=-\sum_{j=1}^{N-1}q_{j}$). It is shown that in the case when $\varrho_{1},\dots,\varrho_{N}\neq 0$, the coefficients $b_{i,j}$ from the relation (1.10) form an invertible matrix, which in turn provides an $L^{2}$ bound on $\nabla q_{i}$. Similar approach, with a different definition of the fluxes, was considered for example in [9, 11].
@@ 2 A priori estimates
$2\sum_{i=1}^{N}\frac{\gamma_{i}}{m_{i}}\int_{\Omega}\varrho_{i}^{\gamma_{i}-1}\omega_{i}\;\mathrm{d}x\geq-2\frac{\gamma_{3}}{m_{3}}\int_{\Omega}\varrho_{3}^{\gamma_{3}-1}\omega_{3}\;\mathrm{d}x.$
@@
[…]
@@ 3.1 The Galerkin approximation
The equation (3.3) admits a global weak solution in $0,T]\times\Omega$.
@@ 3.3 Boundedness of $\varrho$
$2\sum_{i=1}^{N}\frac{1}{m_{i}}\int_{\Omega}\varrho_{i}\omega_{i}^{\delta}\;\mathrm{d}x\geq-2\frac{1}{m_{3}}\int_{\Omega}\varrho_{3}\omega_{3}^{\delta}\;\mathrm{d}x.$
@@
By performing elementary operations on the matrix $B=(b_{i,j})_{i,j=1,\dots,N-1}$, it is easy to see that $\det B=\frac{1}{\varrho}\varrho_{1}\dots\varrho_{N}$. Indeed, adding the verses $2,\dots,N-1$ to the first one and dividing it by $\frac{\varrho_{N}}{\varrho}$, we get the matrix
@@
$\det B=\frac{\varrho_{N}}{\varrho}\cdot\varrho_{1}\dots\varrho_{N-1}\text{ \ \ as we claimed.}$
@@
$\limsup_{k}\int_{0}^{T}\frac{1}{\|K_{h}\|_{1}}\iint_{\Omega\times\Omega}K_{h}(x-y)|\varrho_{k}(t,x)-\varrho_{k}(t,y)|^{p}\;\mathrm{d}x\mathrm{d}y\mathrm{d}t\to 0\quad\text{as}\quad h\to 0,$
then $\{\varrho_{k}\}_{k\in\mathbb{N}}$ is compact in $L^{p}((0,T)\times\Omega)$. Conversely, if $\{\varrho_{k}\}_{k\in\mathbb{N}}$ is compact in $L^{p}((0,T)\times\Omega)$, then the above quantity converges to 0.
@@
The last term has a good sign, so we just estimate it by 0. The first term is estimated in the same way as in [7]. By the definition of $K_{h}$, we have
@@
Denoting by EdEd an $H^{1}(\mathbb{R}^{3})$ extension of $d$, we have
@@
applied to $M|\nabla u^{x}|$ and $\int_{\Omega}\overline{K}_{h}(x-y)(\varrho^{x}-\varrho^{y})^{2}\mathrm{d}y$ extended by 0 outside $\Omega$. We get
@@ 5 The general diffusing/non-diffusing case
and computing $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{R}_{h}$, it is clear that for the non-diffusing components the equivalents of the terms $A_{1}$, $A_{2}$, $A_{4}$ and $A_{5}$ can be dealt with in the same manner, whereas $A_{3}$ is just equal to 0. In consequence, the computations from Section 4 give the strong convergence of densities for both diffusing and non-diffusing components.